# Teaching an Old Dog New Tricks: Verifiable FHE Using Commodity Hardware

### v1

@@ Abstract.
Argos requires no dedicated hardware extensions and is supported on commodity processors from 2008 onward. Our prototype implementation introduces 6% overhead to the FHE evaluation, and 8% for more complex protocols. In particular, we show that Argos can be used for real-world applications of FHE, such as private information retrieval (PIR) and private set intersection (PSI), where providing verifiability is imperative. By demonstrating how to combine cryptography with trusted hardware, Argos paves the way for widespread deployment of FHE-based protocols beyond the semi-honest setting, without the overhead of cryptographic proofs.
@@ 1. Introduction
At the forefront of these issues, the fact that FHE-schemes are not secure when considering a malicious evaluator, or that existing FHE schemes lack notions of integrity, are significant limitations for real-world deployment (Atapoor et al., 2024; Viand et al., 2023; Chase et al., 2017). First, if a malicious server is able to supply malformed ciphertext for a client to decrypt, it can mount key recovery attacks (Guo et al., 2024; Chenal and Tang, 2015) and break all security (if the attacker recovers the private key, it can now decrypt the client’s private requests). Second, if FHE enables delegation of computation on confidential data to an untrusted party, existing constructions cannot help the user verify that the correct function was evaluated. In other words, FHE schemes always assumes honest-but-curious attackers. In the case of smart contracts, for instance, a malicious evaluator can completely change the functionality of the contract, with the option to even reveal confidential inputs and violate privacy. A straightforward solution is to add a layer of verifiable computation on top of the FHE scheme. Unfortunately, existing solutions suffer from major limitations. Cryptographic proofs incur an overhead of 4 to 6 orders of magnitude on top of FHE (Viand et al., 2023). Another solution is to replicate the evaluation across several non-colluding parties, but distributed-trust setups are often impractical in the real world (Lindell et al., 2023).
@@
Modern platforms require dedicated hardware extensions and are currently only available on server-grade processors, limiting adoption.
@@
This Work. We present Argos, the first enclave platform specifically designed to build maliciously-secure and verifiable FHE. As observed in previous work (Kumar et al., 2020; Natarajan et al., 2021; Viand et al., 2023; Tramer et al., 2017), for use cases where all application data is encrypted, such as in FHE, the only secrets exposed on the server are those used for attestation. We go further and show how this is a unique opportunity to rethink TEE architecture and design integrity-only enclaves that focus on data and code integrity without offering any confidentiality guarantee. Lessening the requirements for enclaves gives us the opportunity to 1) completely eliminate microarchitectural side channels by secluding all secret key material in a physically separated coprocessor (TPM), 2) use a hypervisor-based TEE architecture that offers better performance and hardware compatibility, and 3) build a simplified attestation scheme that makes it possible to easily prove the security of our construction and amortize the cost of relying on a slow coprocessor for key operations.
@@
Eliminating Microarchitectural Side Channels. An important observation we make is that side-channel attacks are “read-only” gadgets and can be eliminated simply by excluding any secrets from the CPU and the memory hierarchy (see fig. 1). Software-mounted side channels require an attacker program to share hardware resources with the victim programs in order to extract secrets. If all secrets are secluded on a separate chip, such as a physical trusted platform module (TPM) or a secure coprocessor (Corporation, 2022; Apple, 2022; Johnson et al., 2018), no attacker program can ever share microarchitectural resources with the cryptographic algorithm and extract secrets. Leveraging this insight, we design Argos to store all attestation secrets inside a discrete TPM and delegate all sensitive operations outside of the CPU. This simple principle makes Argos secure by construction against side-channels. The only (side) channel left between the secluded cryptographic keys and a potential attacker is completion time (Moghimi et al., 2020) which is addressed by using constant time cryptography in the TPM (see our threat model in section 2).
@@
Rehabilitating Hypervisor-based TEEs. Argos relies on a hypervisor-based TEE design. A security monitor runs with hypervisor privilege and provides code integrity and isolation to enclave programs. Such TEEs were introduced by TrustVisor (McCune et al., 2010) and are still widely used by cloud vendors to implement confidential computing (Apple, 2024). These architectures are usually considered less secure than dedicated hardware platforms such as Intel SGX or AMD-SEV as they do not encrypt main memory and offer no security against Coldboot (Halderman et al., 2009), one of the most practical physical attacks (see table 7). In our context, this is not an issue as no secrets are ever exposed to main memory. As a result, despite our strong threat model, Agros can adapt this architecture to bring compatibility with most commodity hardware. In terms of performance, existing platforms typically rely on virtual TPMs (Raj et al., 2016)—one per execution environment—that are endorsed by the root TPM but executed on CPU (Azure, 2024; AWS, 2024b). While this approach avoids the performance bottleneck of a single physical TPM (McCune et al., 2010), it makes them vulnerable to microarchitectural side-channels. In contrast, Argos prefers a secure-by-design approach, centered on the discrete TPM.
@@
Performance Optimization With a Discrete Chip. Relying on a single discrete TPM can have a significant impact on performance (McCune et al., 2008). Nevertheless, we can still virtualize most TPM resources in the security monitor as measurement (e.g., hashing) does not require any secret manipulation. Our use case also excludes the use of the TPM “sealed” secret storage. That means we only use the TPM as a “signing oracle.” This can still be expensive, especially when considering interactive cryptographic protocols that require repeated and attested message exchanges.
@@
Simple and Efficient Attestation Scheme. We show that for our FHE applications, we can build a simple attestation scheme that only ever requires one TPM signature. Unlike in other enclave platforms, the remote user does not need to see any attestation proof before sending its sensitive data, as it will always stay encrypted. This also applies to intermediate protocol messages that do not need to be attested but simply added to a transcript whose hash is extended in the security monitor. At the end of the FHE computation (or more precisely, before any FHE ciphertext needs to be decrypted), the transcript’s hash will be signed using the TPM. We show how our simplified transcript-based attestation can be model as a proof system, and show it is sufficient to achieve malicious security needed for verifiable FHE. We also show how Argos can be extended to support batched verifiable FHE. As a result, the overhead of using a physical TPM becomes a fixed cost, and performance becomes similar to that of a virtual TPM, while enforcing security-with-abort in the presence of a malicious attacker (Goldreich et al., 2019).
@@
Implementation And Evaluation. We implement and evaluate Argos on real hardware. Our prototype runs on Intel x86 platforms, but Argos is easily adaptable to other vendors and architectures (e.g., AMD X86, ARM or RISC-V) and compatible with most commodity machines past 2008. To guarantee the TCB integrity at boot-time, we use commodity hardware root-of-trust technologies. Our prototype uses a TPM, but our architecture is adaptable to other hardware roots of trust and secure coprocessors such as the Apple Secure Enclave (Apple, 2022) or Open Titan (Johnson et al., 2018). Our security monitor is an open source fork of Tyche (Castes et al., 2023), modified to support our attestation scheme and remove side-channel protections, but our approach is also compatible with other micro-hypervisors like seL4 (Klein et al., 2009). We implement a custom minimal runtime for FHE and the SEAL library (Chen et al., 2017a), emulating system calls and providing randomness through a hardware random number generator (i.e., RDRAND instructions) that is not under OS control. For more complex applications that require a broader class of system calls, Gramine (Tsai et al., 2017) can be used for better compatibility, at the price of a larger TCB and some performance overhead (5x slower startup time, for example). Argos has a limited attack surface with a minimal trusted code base (TCB) of less than 18KLOC, plus the target FHE application (KLOC). Our evaluation shows that Argos is 80 times faster than previous work leveraging Intel SGX for FHE integrity (Viand et al., 2023) with a minimal average performance overhead of 3% for FHE evaluation. We show that Argos can be used to implement more complex protocols such as attested PIR and attested PSI with performance overheads under 8% and without the significant offline communication costs incurred by cryptographic solutions.
@@
Argos, the first integrity-only enclave platform designed to build maliciously-secure verifiable FHE, which:
can be used to build fully malicious and authenticated PSI and PIR schemes;
is secure by construction against microarchitectural side channels and transient execution attacks;
requires no specialized hardware and is compatible with commodity processors from 2008 onward; and
only incurs 3% average performance overhead for FHE computation, less than 8% performance overhead for complex protocols, and virtually no communication overhead;
@@ 2. Threat model
Our goal is to ensure integrity of the data and computation running inside our enclave environment. Most importantly, we do not protect confidentiality of our programs nor do we protect them against denial of service. We assume a strong adversary collocated on the server that has compromised the majority of the software stack including the OS, and can mount any microarchitectural side-channel and transient execution attacks. Because Argos is secure against all software-mounted side channels that underlie known transient execution attacks, we will refer to all these attacks as “microarchitectural side channels” throughout the rest of the paper. Our trusted computing base (TCB) consists only of a security monitor (18KLOC) and our application, both assumed to be bug-free. We assume that the hardware is functionally correct and that the TPM’s cryptography is properly implemented with constant-time programming. Rowhammer and fault injection attacks are not as practical and are considered out of scope. We also protect against the majority of physical attacks (cold boot, BIOS tampering, physical side channels) on the main processor and DRAM, but for the rest of this paper, we will focus on software-mounted attacks and consider them out of scope. A detailed discussion on the remaining attack surface can be found in section 9.7.
@@ 3.1. FHE Schemes And Client-Server Setup
FHE schemes must enforce the following properties:
Correctness. An scheme is correct if any honest computation will decrypt to the expected result (i.e. with high probability).
@@
In a honest-but-curious or semi-honest setting, at attacker is allowed to observe intermediate messages and passively infer information in order to break security but cannot deviate from the agreed-upon protocol. In a malicious setting, the attacker is much more powerful and allowed to deviate arbitrarily from the protocol, e.g., to lie to the victim or temper with messages. The semi-honest setting is often used in theory but has limited viability in a deployment context where assuming a malicious attacker is more realistic.
@@ 3.3. Why Does FHE Need Verifiability?
Most state-of-the-art FHE schemes are insecure to use in real-world settings (Gentry, 2009; Chillotti et al., 2020; Ducas and Micciancio, 2015; Chen et al., 2017a). Because these schemes are not CCA-secure (Bellare et al., 1998), if an attacker is able to supply malformed ciphertexts and observe decryption outcomes (for example, by observing client behavior), it can mount key recovery attacks and completely compromise privacy (Guo et al., 2024; Fauzi et al., 2022; Chaturvedi et al., 2022; Chenal and Tang, 2015). Beyond the “circuit-level” security, the absence of integrity for FHE schemes also has implications for FHE-based applications. A malicious server can arbitrarily modify the circuit they evaluate on the encrypted data. Attacks on correctness are an obvious issue, but these can also translate to attacks on privacy at the application level. Let us take the example of private contact discovery. Alice wants to discover who among her contacts is using a private messaging service (moxie0, 2017). Alice encrypts her address book, which contains all of her contacts’ private information, and sends it to the server. The server now has access to Alice’s encrypted set of contacts. If malicious, it could simply return the entire set to Alice. This could mislead Alice into thinking that one of her friends (Bob) uses the messaging service. Consequently, Alice contacts Bob via the service, leaking his phone number. This would violate Bob’s privacy and defeat the use of private contact discovery.
@@ 3.4. Why do we need a new TEE platform?
Trusted execution environments (TEEs) provide hardware-based isolation for confidential programs while minimizing the trusted code base (see fig. 1). The existing TEE landscape spans from platforms that protect small “enclave” programs (AWS, 2024a; Pinto and Santos, 2019; Costan and Devadas, 2016) to those that implement confidential computing for entire virtual machines (Li et al., 2022; Cheng et al., 2024; SEV-SNP, 2020; Apple, 2024). However, the discovery of numerous microarchitectural side-channel and transient execution attacks has severely compromised their confidentiality guarantees. Side channel attacks exploit shared microarchitectural structures such as memory caches (Götzfried et al., 2017; Brasser et al., 2017), translation look-aside buffers (TLB) (Gras et al., 2018), branch predictors (Evtyushkin et al., 2018; Acıiçmez et al., 2006), and DRAM controllers (Wang et al., 2014) to enable information leakage between security domains.
TEEs also provide integrity guarantees through remote attestation, allowing a remote client to verify the authenticity of the hardware and the initial state of the execution environment. However, even these attestation mechanisms have proven vulnerable to side-channel attacks that can extract secret keys used by platforms to sign attestation reports (van Schaik et al., 2020; Chen et al., 2019; Ryan, 2019), leading the cryptographic community to lose faith in TEEs. On the other hand, discrete trusted platform modules (TPMs) and associated dynamic root of trust (DRoT) have long served as dedicated hardware for platform integrity and secure storage against software-based attackers. However, their discrete nature also creates a severe performance bottleneck when attesting multiple security domains (Flicker (McCune et al., 2008) incurs 2-3 order of magnitude overheads). Although TPM virtualization (Raj et al., 2016; AWS, 2024b; Azure, 2024; McCune et al., 2010) can address these performance issues, in turn they expose sensitive key material to microarchitectural side channels (see section 10 for more detailed comparison of Argos and other platforms).

### v2

@@ Abstract.
Argos requires no dedicated hardware extensions and is supported on commodity processors from 2008 onward. Our prototype implementation introduces 3% overhead for FHE evaluation, and 8% for more complex protocols. In particular, we show that Argos can be used for real-world applications of FHE, such as private information retrieval (PIR) and private set intersection (PSI), where providing verifiability is imperative. By demonstrating how to combine cryptography with trusted hardware, Argos paves the way for widespread deployment of FHE-based protocols beyond the semi-honest setting, without the overhead of cryptographic proofs.
@@ 1. Introduction
At the forefront of these issues, the fact that FHE-schemes are not secure when considering a malicious evaluator, or that existing FHE schemes lack notions of integrity, are significant limitations for real-world deployment (Atapoor et al., 2024; Viand et al., 2023; Chase et al., 2017). First, if a malicious server is able to supply malformed ciphertext for a client to decrypt, it can mount key recovery attacks (Guo et al., 2024; Chenal and Tang, 2015) and break all security (if the attacker recovers the private key, it can now decrypt the client’s private requests). Second, if FHE enables delegation of computation on confidential data to an untrusted party, existing constructions cannot help the user verify that the correct function was evaluated. In other words, FHE schemes always assume honest-but-curious attackers. In the case of smart contracts, for instance, a malicious evaluator can completely change the functionality of the contract, with the option to even reveal confidential inputs and violate privacy. A straightforward solution is to add a layer of verifiable computation on top of the FHE scheme. Unfortunately, existing solutions suffer from major limitations. Cryptographic proofs incur an overhead of 4 to 6 orders of magnitude on top of FHE (Viand et al., 2023). Another solution is to replicate the evaluation across several non-colluding parties, but distributed-trust setups are often impractical in the real world (Lindell et al., 2023).
@@
Modern platforms require dedicated hardware and are only available on server-grade processors, limiting adoption.
@@
This Work. We present Argos, the first enclave platform specifically designed to build maliciously-secure and verifiable FHE. As observed in previous work (Kumar et al., 2020; Natarajan et al., 2021; Viand et al., 2023; Tramer et al., 2017), for use cases where all application data is encrypted, such as in FHE, the only secrets exposed on the server are those used for attestation. We go further and show how this is a unique opportunity to rethink TEE architecture and design integrity-only enclaves that focus on data and code integrity without offering any confidentiality guarantee. Lessening the requirements for enclaves gives us the opportunity to 1) completely eliminate microarchitectural side channels by secluding all secret key material in a physically-separated coprocessor (such as a trusted platform module or TPM), 2) use a hypervisor-based TEE architecture that offers better performance and hardware compatibility, and 3) build a simplified attestation scheme that makes it possible to easily prove the security of our construction and amortize the cost of relying on a slow coprocessor for key operations.
@@
Eliminating Microarchitectural Side Channels. An important observation we make is that side-channel attacks are “read-only” gadgets and can be eliminated simply by excluding any secrets from the CPU and the memory hierarchy (see Figure 1). Software-mounted side channels require an attacker program to share hardware resources with the victim programs in order to extract secrets. If all secrets are secluded on a separate chip, such as a physical trusted platform module (TPM) or a secure coprocessor (Corporation, 2022; Apple, 2022; Johnson et al., 2018), no attacker program can ever share microarchitectural resources with the cryptographic algorithm and extract secrets. Leveraging this insight, we design Argos to store all attestation secrets inside a discrete TPM and delegate all sensitive operations to outside of the CPU. This simple principle makes Argos secure by construction against side-channels. The only (side) channel left between the secluded cryptographic keys and a potential attacker is completion time (Moghimi et al., 2020) which is addressed through using constant-time cryptography in the TPM (see our threat model in Section 2).
@@
Rehabilitating Hypervisor-based TEEs. Argos relies on a hypervisor-based TEE design. A security monitor runs with hypervisor privilege and provides code integrity and isolation to enclave programs. Such TEEs were introduced by TrustVisor (McCune et al., 2010) and are still widely used by cloud vendors to implement confidential computing (Apple, 2024a). These architectures are usually considered less secure than dedicated hardware platforms such as Intel SGX or AMD-SEV as they do not encrypt main memory and offer no security against Coldboot (Halderman et al., 2009), one of the most practical physical attacks (see Table 7). In our context, this is not an issue as no secrets are ever exposed to main memory. As a result, despite our strong threat model, Argos can adapt this architecture to bring compatibility with most commodity hardware. In terms of performance, existing platforms typically rely on virtual TPMs (Raj et al., 2016)—one per execution environment—that are endorsed by the root TPM but executed on CPU (Azure, 2024; AWS, 2024b). While this approach avoids the performance bottleneck of a single physical TPM (McCune et al., 2010), it makes them vulnerable to microarchitectural side-channels. In contrast, Argos prefers a secure-by-design approach, centered on the discrete TPM.
@@
Performance Optimization With a Discrete Chip. Relying on a single discrete TPM can have a significant impact on performance (McCune et al., 2008). Nevertheless, we can still virtualize most TPM resources in the security monitor, trusted code that runs at hypervisor privilege level, as measurement (e.g., hashing) does not require any secret manipulation. Our use case also excludes the use of the TPM “sealed” secret storage. That means we only use the TPM as a “signing oracle”. This can still be expensive, especially when considering interactive cryptographic protocols that require repeated and attested message exchanges.
@@
Simple and Efficient Attestation Scheme. We show that for our FHE applications, we can build a simple attestation scheme that only ever requires one TPM signature. Unlike in other enclave platforms, the remote user does not need to see any attestation proof before sending its sensitive data, as it will always stay encrypted. This also applies to intermediate protocol messages that do not need to be attested but simply added to a transcript whose hash is extended in the security monitor. At the end of the FHE computation (or more precisely, before any FHE ciphertext needs to be decrypted), the transcript’s hash will be signed using the TPM. We show how our simplified transcript-based attestation can be model as a proof system, and show that it is sufficient to achieve malicious security needed for verifiable FHE. We also show how Argos can be extended to support batched verifiable FHE. As a result, the overhead of using a physical TPM becomes a fixed cost, and performance becomes similar to that of a virtual TPM, while enforcing security-with-abort in the presence of a malicious attacker (Goldreich et al., 2019).
@@
Implementation and Evaluation. We implement and evaluate Argos on real hardware111https://github.com/mit-enclaves/argos. Our prototype runs on Intel x86 platforms, but Argos is easily adaptable to other vendors and architectures (e.g., AMD X86, ARM or RISC-V) and compatible with most commodity machines past 2008. To guarantee the TCB integrity at boot-time, we use commodity hardware root-of-trust technologies. Our prototype uses a TPM, but our architecture is adaptable to other hardware roots of trust and secure coprocessors such as the Apple Secure Enclave (Apple, 2022) or Open Titan (Johnson et al., 2018). Our security monitor is an open source fork of Tyche (Castes et al., 2023), modified to support our attestation scheme and remove side-channel protections, but our approach is also compatible with other micro-hypervisors like seL4 (Klein et al., 2009). We implement a custom minimal runtime for FHE and the SEAL library (Chen et al., 2017a), emulating system calls and providing randomness through a hardware random number generator (i.e., RDRAND instructions) that is not under OS control. For more complex applications that require a broader class of system calls, Gramine (Tsai et al., 2017) can be used for better compatibility, at the price of a larger TCB and some performance overhead (5 slower startup time, for example). Argos has a limited attack surface with a minimal trusted code base (TCB) of less than 18KLOC, plus the target FHE application (KLOC). Our evaluation shows that Argos is 80 times faster than previous work leveraging Intel SGX for FHE integrity (Viand et al., 2023) with a minimal average performance overhead of 3% for FHE evaluation. We show that Argos can be used to implement more complex protocols such as attested PIR and attested PSI with performance overheads under 8% and without the significant offline communication costs incurred by cryptographic solutions.
Contribution. Argos is the first TEE-based platform to enable maliciously-secure verifiable FHE while being secure against all known microarchitectural side channels, all at minimal overheads. Prior to this work, it was not obvious that TEEs in commodity hardware could achieve such strong security guarantees. Existing TEEs suffer from insecurity due to the fact that their remote attestation mechanisms are vulnerable to microarchitectural side channels (see Section 9.7). As a result, TEEs have thus far remained undesirable for cryptographic applications such as verifiable FHE. Argos is carefully designed such that only the TPM contains unencrypted secrets. The TPM uses a simple microarchitecture and is microarchitecturally isolated from the CPU, blocking all known microarchitectural side channels. This means that we can leverage the TPM as a ”signing oracle” to design a simple, efficient, and secure remote attestation scheme. Furthermore, this simple solution addresses an important problem: the efficient deployment of FHE in real-world scenarios with malicious security.
@@
Argos, the first integrity-only enclave platform designed to build maliciously-secure verifiable FHE;
Argos can be used to build fully malicious and authenticated PSI and PIR schemes;
Argos is secure by construction against microarchitectural side channels and transient execution attacks;
Argos requires no specialized hardware and is compatible with commodity processors from 2008 onward; and
Argos only incurs 3% average performance overhead for FHE computation, less than 8% performance overhead for complex protocols, and virtually no communication overhead;
@@ 2. Threat model
Our goal is to ensure integrity of the data and computation running inside our enclave environment. Most importantly, we do not protect confidentiality of our programs, nor do we protect them against denial of service. We assume a strong adversary collocated on the server that has compromised the majority of the software stack including the OS, and can mount any microarchitectural side-channel and transient execution attacks. Because Argos is secure against all software-mounted side channels that underlie known transient execution attacks, we will refer to all these attacks as “microarchitectural side channels” throughout the rest of the paper. Our trusted computing base (TCB) consists only of a security monitor (18 KLOC) and our application, both assumed to be bug-free. We assume that the hardware is functionally correct and that the TPM’s cryptography is properly implemented with constant-time programming. Rowhammer and fault injection attacks are not as practical and are considered out of scope. We also protect against the majority of physical attacks (cold boot, BIOS tampering, physical side channels) on the main processor and DRAM, but for the rest of this paper, we will focus on software-mounted attacks and consider physical attacks out of scope. A primer on microarchitectural and physical attacks can be found in Appendix B. A detailed discussion on the remaining attack surface can be found in Section 9.7.
Evolution of the attack surface on TEE platforms. Are presented: baseline, dynamic root of trust, hypervisor-based, SGX, and Argos. Argos has a significantly reduced attack surface.
@@ 3.1. FHE Schemes And Client-Server Setup
FHE schemes must enforce the following properties (see Appendix C.1 for formal definitions):
Correctness. A scheme is correct if any honest computation will decrypt to the expected result (i.e., with high probability).
@@
In a honest-but-curious or semi-honest setting, an attacker is allowed to observe intermediate messages and passively infer information in order to break security, but cannot deviate from the agreed-upon protocol. In a malicious setting, the attacker is much more powerful and allowed to deviate arbitrarily from the protocol, e.g., to lie to the victim or tamper with messages. The semi-honest setting is often used in theory, but has limited viability in a deployment context where assuming a malicious attacker is more realistic.
@@ 3.3. Why Does FHE Need Verifiability?
Most state-of-the-art FHE schemes are insecure to use in real-world settings (Gentry, 2009; Chillotti et al., 2020; Ducas and Micciancio, 2015; Chen et al., 2017a) where malicious security is generally required. Because these schemes are not CCA-secure (Bellare et al., 1998), if an attacker is able to supply malformed ciphertexts and observe decryption outcomes (for example, by observing client behavior), it can mount key recovery attacks and completely compromise privacy (Guo et al., 2024; Fauzi et al., 2022; Chaturvedi et al., 2022; Chenal and Tang, 2015). Beyond the “circuit-level” security, the absence of integrity for FHE schemes also has implications for FHE-based applications. A malicious server can arbitrarily modify the circuit they evaluate on the encrypted data. Attacks on correctness are an obvious issue, but these can also translate to attacks on privacy at the application level. Let us take the example of private contact discovery. Alice wants to discover who among her contacts is using a private messaging service (moxie0, 2017). Alice encrypts her address book, which contains all of her contacts’ private information, and sends it to the server. The server now has access to Alice’s encrypted set of contacts. If malicious, it could simply return the entire set to Alice. This could mislead Alice into thinking that one of her friends (Bob) uses the messaging service. Consequently, Alice contacts Bob via the service, leaking his phone number. This would violate Bob’s privacy and defeat the use of private contact discovery.
@@ 3.4. Why Do We Need a New TEE Platform?
Trusted execution environments (TEEs) provide hardware-based isolation for confidential programs while minimizing the trusted code base (see Figure 1). The existing TEE landscape spans from platforms that protect small “enclave” programs (AWS, 2024a; Pinto and Santos, 2019; Costan and Devadas, 2016) to those that implement confidential computing for entire virtual machines (Li et al., 2022; Cheng et al., 2024; SEV-SNP, 2020; Apple, 2024a). However, the discovery of numerous microarchitectural side channel and transient execution attacks has severely compromised their confidentiality guarantees. Side channel attacks exploit shared microarchitectural structures such as memory caches (Götzfried et al., 2017; Brasser et al., 2017), translation look-aside buffers (TLB) (Gras et al., 2018), branch predictors (Evtyushkin et al., 2018; Acıiçmez et al., 2006), and DRAM controllers (Wang et al., 2014) to enable information leakage between security domains.
TEEs also provide integrity guarantees through remote attestation, allowing a remote client to verify the authenticity of the hardware and the initial state of the execution environment. However, even these attestation mechanisms have proven vulnerable to side-channel attacks that can extract secret keys used by platforms to sign attestation reports (van Schaik et al., 2020; Chen et al., 2019; Ryan, 2019), leading the cryptographic community to lose faith in TEEs. On the other hand, discrete trusted platform modules (TPMs) and associated dynamic root of trust (DRoT) have long served as dedicated hardware for platform integrity and secure storage against software-based attackers. However, their discrete nature also creates a severe performance bottleneck when attesting multiple security domains (Flicker (McCune et al., 2008) incurs 2-3 orders of magnitude overhead). Although TPM virtualization (Raj et al., 2016; AWS, 2024b; Azure, 2024; McCune et al., 2010) can address these performance issues, in turn, they expose sensitive key material to microarchitectural side channels (see Section 10 for a more detailed comparison of Argos and other platforms).